OpenStudy (anonymous):

Please help me with this question! I don't know what to do! The diameter of circle center Z is 5in. What is its area in terms of pi?

OpenStudy (whpalmer4):

\[d = 2r\]\[r=\frac{d}2\]\[ A = \pi r^2 =\pi(\frac{d}2)^2 =\frac{\pi d^2}{4}\] Plug in 5 inches for the value of \(d\)

OpenStudy (anonymous):

I got 25. Is that right?

OpenStudy (whpalmer4):

How do you figure that? Can you show me your work?

OpenStudy (whpalmer4):

Don't bump the post, answer my question...

OpenStudy (whpalmer4):

If you show me where you made a mistake, we can clear up the misunderstanding. If you just get someone else to give you the answer, you'll have the answer, but no understanding.

OpenStudy (whpalmer4):

For starters, if someone says "give me <something> in terms of <something else>" your answer needs to have <something else> in it...

OpenStudy (anonymous):

Sorry. That was an accident. I'm trying to redo it because it make sense to me. This time I got 2.5^2

OpenStudy (whpalmer4):

I still don't see \(\pi\) in there anywhere...and the number isn't right either. can you please show me your work, as I asked?

OpenStudy (whpalmer4):

Sorry, the number will be right. Need to clean my screen, it looked like 2.6^2!

OpenStudy (anonymous):

Okay. So I pIuggd in 5 into 'd' as you've asked. 5=2r so then I did this: r=5/2=2.5 Then I got lost at the last part.

OpenStudy (whpalmer4):

Okay, I then gave you a formula that used d in place of r, didn't I? You can plug 5 directly into that: \[A = \frac{\pi d^2}{4} = \frac{\pi(5)^2}{4} =\]?

OpenStudy (anonymous):

So (5)/4 means 25/5 right? I just want to make sure I'm doing this right.

OpenStudy (whpalmer4):

What is \((5)^2\)?

OpenStudy (whpalmer4):

25, right? so \[\frac{(5)^2}{4} = \frac{25}{4}\]

OpenStudy (whpalmer4):

So the answer is \[\frac{25\pi}{4}\]or if you want a decimal expression you could write \[6.25\pi\]but most people would prefer the fraction, I think.

OpenStudy (whpalmer4):

Depends on what you're going to do with the result — if you are computing a number, then the decimal might be preferable. If you're doing some more algebra, the fraction might be better so you could spot cancellation.

OpenStudy (anonymous):

Okay. I got it right then. Thanks so much and sorry for all the frustration.

OpenStudy (whpalmer4):

You're welcome. Sorry for misreading your result!